Exercise 11-3 (Part Level Submission)

Monty Company purchased a new plant asset on April 1, 2017, at a cost of $774,990. It was estimated to have a service life of 20 years and a salvage value of $65,400. Monty’s accounting period is the calendar year.

Collapse question part

(a)

Incorrect answer. Your answer is incorrect. Try again.
Compute the depreciation for this asset for 2017 and 2018 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.)
Depreciation for 2017 $

Entry field with incorrect answer now contains modified data

Depreciation for 2018 $

Entry field with incorrect answer now contains modified data
Answer #1

Sum of year digit = 210

Depreciation for 2017 = (774990-65400)*20/210*9/12 = 50685

Depreciation for 2018 = 709590*20/210*3/12+709590*19/210*9/12 = 65046